# Asymptotic stability equals exponential stability—while you twist your eyes

### v1

@@ 1.3   Stability
Let $\gamma\in\mathcal{K}_{\infty}$ be smooth on $(0,+\infty)$ and such that $\gamma(s)/\gamma^{\prime}(s)\geq s$ (e.g., $s\mapsto\gamma(s)=s^{(1/2)}$). Now consider the function $x\mapsto V_{\gamma}(x):=\gamma(\|x\|_{2})$ on $\mathbb{R}^{n}$, which is $C^{\infty}$ on $\mathbb{R}^{n}\setminus\{0\}$. Next, construct the vector field
$\dot{x}=X_{3}(x):=\begin{cases}-\nabla V_{\gamma}(x)\quad&x\neq 0\\ 0\quad&x=0\end{cases}.$ (1.6)
We cannot immediately appeal to Filippov’s framework as $X_{3}$ is not necessarily a bounded operator. However, since $-\nabla V_{\gamma}(x)=-\gamma^{\prime}(\|x\|_{2})x/\|x\|_{2}$ we can study solutions to (1.6) directly. To that end, decompose $x\in\mathbb{R}^{n}\setminus\{0\}$ as $x=\|x\|_{2}\cdot x/\|x\|_{2}=:r\cdot u$. It readily follows that $\dot{r}=-\gamma^{\prime}(r)$ while $\dot{u}=0$. In general $\gamma^{\prime}(s)>0$ $\forall s\geq 0$, need not be true, but suppose this is true for our choice of $\gamma$, e.g., pick again $s\mapsto\gamma(s)=s^{1/2}$. Now, suppose that $1/\gamma^{\prime}(s)$ is continuous and of class $\mathcal{K}$ on $\mathbb{R}_{\geq 0}$, then, we can define $\Gamma(r):=\int^{r}_{0}1/\gamma^{\prime}(\rho)\mathrm{d}\rho$, which is now of class $\mathcal{K}_{\infty}$ and hence invertible on $\mathbb{R}_{\geq 0}$. Under the aforementioned assumptions, we can define a semiflow corresponding to $X_{3}$ via (e.g., derived via the inverse function theorem):
$(t,x)\mapsto\varphi_{3}^{t}(x):=\begin{cases}\min\{0,\Gamma^{-1}(\Gamma(\|x\|_{2})-t)\}x/\|x\|_{2}\quad&\|x\|_{2}\geq\Gamma^{-1}(t)\\ 0\quad&\mathrm{else}\end{cases}.$
We emphasize that $\lim_{x\to 0}\varphi_{3}^{t}(x)=0$ for any $t\geq 0$, which follows for $t=0$ from $\varphi^{0}_{3}=\mathrm{id}_{\mathbb{R}^{n}}$, whereas for $t>0$ we have that $\min\{0,\Gamma^{-1}(\Gamma(\|x\|_{2})-t)\}=0$ for all $\|x\|_{2}\leq\Gamma^{-1}(t)$. Note, the latter is inherent to the definition of $\Gamma$. $\circ$
@@ 2   Further comments on related work
Recently, we showed that when the origin $0\in\mathbb{R}^{n}$ is GAS under a continuous vector field $X$, and if this can be asserted using a $C^{1}$ convex Lyapunov function $V$, then, $X$ is straight-line homotopic to $-\partial_{x}$, such that the origin remains GAS along the homotopy [ref:JongeneelSchwan2024TAC]. This is a convenient result, but clearly not a general one.
Earlier, the following homotopy on the level of vector fields, reminiscent of Alexanders’s trick, appeared in several works (e.g., see [ref:ryan1994brockett, p. 1603], [ref:sontag2013mathematical, Thm. 21], [ref:coron2007control, p. 291] and [ref:jongeneel2023topological, Ex. 3.4]):
@@
As for any $t\geq 0$ the only fixed point of $\varphi^{t}(\cdot;X)$ is 00, it follows that for $n=1$, the homotopy (2.1) preserves stability (globally), simply because the sign of $x\mapsto H(s,x)$ cannot flip for otherwise, $\varphi^{s/(1-s)}(x;X)=x$ must hold for some $s\in(0,1)$ and $x\neq 0$. $\circ$
Example 2.2 (Stability preserving homotopies for linear dynamical systems).
Let $0\in\mathbb{R}^{n}$ be GAS under $\dot{x}=X(x):=Ax$, for $A=:TJT^{-1}$ the Jordan form decomposition of $A$. Then it readily follows that $\frac{1}{s}(\varphi^{s/(1-s)}(x;X)-x)=\frac{1}{s}T(e^{s/(1-s)J}-I_{n})T^{-1}x$ such that stability is preserved throughout the homotopy (2.1). $\circ$
Either way, the homotopy (2.1) does already show that there is a homotopy that does not introduce new equilibrium points, along the homotopy. Indeed, this has been generalized recently by Kvalheim to compact attractors on manifolds [ref:kvalheim2022obstructions]. However, all of this is not enough to conclude on the existence of a homotopy that preserves stability.
On the other hand, it is known that there is no reason why stability must be preserved along such a homotopy. In the spirit of [ref:eliashberg2002introduction, Sec. 4.1], consider the following family of linear vector fields on $\mathbb{R}^{2}$:
@@
parametrizing a homotopy from $X(\cdot;0)=\partial_{x}$ to $X(\cdot;1)=-\partial_{x}$ along non-vanishing vector fields on $\mathbb{R}^{n}\setminus\{0\}$. Note that indeed, since $n=2$, the (Hopf) indices [ref:milnor65, p. 32] of these two “qualitatively opposite” vector fields are equal, that is, $\mathrm{ind}_{0}(\partial_{x})=1^{n}=(-1)^{n}=\mathrm{ind}_{0}(-\partial_{x})$. It is precisely this weakness of existing homotopy-invariants that we eventually hope to overcome by studying more restrictive equivalence classes.
@@ 3   Stability preserving homotopies
We first consider coordinates, that is, $0\in\mathbb{R}^{n}$ being (strongly) GAS under some appropriately regular vector field on $\mathbb{R}^{n}$.
@@
The vector field $X$ is a measurable map, possibly set-valued at 00, i.e., $0\mapsto X(0)\in 2^{\mathbb{R}^{n}}$, with $X$ being a bounded operator on $\mathbb{R}^{n}$ and locally Lipschitz away from 00.
It is known that under the conditions of Assumption 3.1, $\mathcal{F}[X]$ is upper semi-continuous and compact, convex valued, which allows for a smooth converse Lyapunov theory, e.g., see the discussions [^3] To add, indeed, the existence of a smooth Lyapunov function for a discontinuous vector field relates to robustness [ref:goebel2012hybrid] (e.g., interpret Definition 1.1 as stability prevailing under all perturbations admissible through the inclusion) and of course to the topology of the space and attractor (e.g., there cannot be a single smooth Lyapunov function on the circle $\mathbb{S}^{1}$ that asserts some point $p^{\star}\in\mathbb{S}^{1}$ is GAS under some (discontinuous) vector field $X$ on $\mathbb{S}^{1}$). in [ref:clarke1998asymptotic].
@@
The proof proceeds in 5 steps. First we show that the time-$t$ map $\varphi^{t}(\cdot;\mathcal{F}[X])$, corresponding to $\dot{x}\in\mathcal{F}[X](x)$, can be homotoped along semiflows to a gradient flow $\varphi^{t}(\cdot;-\nabla V)$, such that along the homotopy the origin remains strongly GAS. In Step (ii), by exploiting symmetry we show that a minor extension of [ref:grune1999asymptotic], allows for showing that for any $\gamma\in\mathcal{K}_{\infty}$ there is a $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ (not just $T\in\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$) such that $V(T^{-1}(x))=\gamma(\|x\|_{2})$. In particular, it follows that $T$ is homotopic to $\mathrm{id}_{\mathbb{R}^{n}}$ along a continuous path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$. Then, in Step (iii), we use this map $T$ to homotope $\varphi^{t}(\cdot;-\nabla V)$ to a semiflow, denoted $\widetilde{\varphi}$, that has $V\circ T^{-1}$ as its Lyapunov function. Again, 00 remains strongly GAS along the path. Next, we show in Step (iv) that $\widetilde{\varphi}^{t}$ can be homotoped to the time-$t$ map $\varphi^{t}(\cdot;-\partial_{x})$. At last, we combine the above and conclude in Step (v).
(i) Since 00 is strongly GAS under $\dot{x}\in\mathcal{F}[X](x)$ and $X$ satisfies Assumption 3.1, there is a Lyapunov pair $V\in C^{\infty}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ and $W\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\};\mathbb{R}_{\geq 0})$, that certifies stability of $0\in\mathbb{R}^{n}$ [ref:clarke1998asymptotic, Thm. 1.3], under any Filippov solution to (1.2). In particular, we can construct the homotopy $H:[0,1]\times\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}^{n}\setminus\{0\}$ through non-vanishing vector fields, defined by $(s,x)\mapsto H(s,x):=(1-s)\mathcal{F}[X](x)-s\nabla V(x)$. Then, as for any $s\in[0,1]$ we have that $(1-s)\mathcal{F}[X]-s\nabla V$ also satisfies Assumption 3.1, plus
$\langle\nabla V(x),H(s,x)\rangle\leq-(1-s)W(x)-s\langle\nabla V(x),\nabla V(x)\rangle,\,\,\forall\,x\in\mathbb{R}^{n}\setminus\{0\},$
we find that 00 is strongly GAS, in the sense of Filippov, under $\dot{x}\in(1-s)\mathcal{F}[X](x)-s\nabla V(x)$ for any fixed $s\in[0,1]$. Now, to construct a homotopy on the level of semiflows, we follow Hartman [hartman2002ordinary, Ch. V, p. 93] and consider the following extended system $\Sigma_{H}$ on $[0,1]\times\mathbb{R}^{n}$:
@@
Observe the following, $0\in\mathbb{R}^{n}$ is strongly GAS under $\dot{x}\in(1-s)\mathcal{F}[X](x)-s\nabla V(x)$ for any fixed $s\in[0,1]$, which we can assert by means of the pair $(V,(1-s)W+s\|\nabla V\|_{2}^{2})$. Therefore, as $V$ has compact sublevel sets [ref:clarke1998asymptotic, (L2)], any Filippov solution to $\Sigma_{H}$ is defined for all $t\geq 0$, e.g., see [ref:khalil2002nonlinear, Thm. 3.3] for the standard ODE case. Now since $(s,x)\mapsto H(s,x)$ is locally Lipschitz away from 00, we can appeal to the Picard-Lindelöf theorem [ref:teschl2012ordinary, Thm. 2.2] and hence, by strong asymptotic stability [ref:clarke1998asymptotic, Def. 2.1] cf. Example 1.3, any solution to $\Sigma_{H}$ is also unique, which allows us to define the time-$t$ map $(s,x)\mapsto\varphi^{t}((s,x);\Sigma_{H})$ for any $t\geq 0$. In its turn, by considering the path $[0,1]\ni s\mapsto\varphi^{t}((s,\cdot);\Sigma_{H})$, we see that this time-$t$ map $\varphi^{t}$ defines a homotopy from the semiflow under $\mathcal{F}[X]$ to the (semi)flow under $-\nabla V$. Specifically, we have that $\varphi^{t}((0,x);\Sigma_{H})=(0,\varphi^{t}(x;\mathcal{F}[X]))$, $\varphi^{t}((1,x);\Sigma_{H})=(1,\varphi^{t}(x;-\nabla V))$ and $\varphi^{t}((s,0);\Sigma_{H})=(s,0)$ $\forall s\in[0,1]$. Now, for any fixed $s\in[0,1]$, overload notation and define the time-$t$ map $\varphi^{t}(\cdot;(s,\Sigma_{H})):=\pi_{2:n+1}\circ\varphi^{t}((s,\cdot);\Sigma_{H})$, for $\pi_{2:n+1}$ the projection on the last $n$ coordinates. It follows that this map checks out as a semiflow, since $\pi_{2:n+1}$ is continuous and for any $x\in\mathbb{R}^{n}$ we have that
@@
(ii) Grüne, Wirth and Sontag showed that for our Lyapunov function $V$ at hand, we can always find a $T\in\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$, with $T(0)=0$, such that $V(T^{-1}(x))=\gamma(\|x\|_{2})$, for some $\gamma\in\mathcal{K}_{\infty}$, smooth on $(0,+\infty)$ [ref:grune1999asymptotic, Prop. 1]. We recall their construction of this map. They define $(t,x)\mapsto\psi(t,x)$ to be the (local) flow with respect to the ODE
@@
It follows that, on the domain of $\psi$, $V(\psi(t,x))=V(x)+t$. Now fix some $c>0$, then they define the map $\pi_{c}:\mathbb{R}^{n}\setminus\{0\}\to V^{-1}(c)$ by $x\mapsto\pi_{c}(x):=\psi(c-V(x),x)$, that is, starting from $x\in\mathbb{R}^{n}\setminus\{0\}$ you flow—either backward or forward—along $\psi$ until you hit $V^{-1}(c)$. Then, due to initial work by Wilson [ref:wilson1967structure], we know that $V^{-1}(c)\simeq_{h}\mathbb{S}^{n-1}$, however, now—that is, after [ref:grune1999asymptotic] was published, the resolution of the Poincaré conjecture [^4] Specifically, Perelman provided the final step $(\mathbb{S}^{3})$ in proving the generalized Poincaré conjecture in $\mathsf{Top}$. For some historical comments, see [ref:stillwell2012poincare]. implies there must be a homeomorphism $S:V^{-1}(c)\to\mathbb{S}^{n-1}$, for any $n\geq 1$. Next, define the map $Q:=S\circ\pi_{c}:\mathbb{R}^{n}\setminus\{0\}\to\mathbb{S}^{n-1}$ and eventually the map $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by
@@
$x\mapsto T(x):=\begin{cases}\gamma^{-1}(V(x))Q(x)\quad&\forall\,x\in\mathbb{R}^{n}\setminus\{0\}\\ 0\quad&\text{otherwise},\end{cases}$
@@
Now, it turns out that the particular choice of the homeomorphism $S:V^{-1}(c)\to\mathbb{S}^{n-1}$ is irrelevant for the construction of $T$ as in [ref:grune1999asymptotic, Prop. 1]. Hence, we simply do the following. If $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, that is, if $T$ preserves orientation, we are done. If not, we can always adjust $T$, for instance, we can compose $S$ with a map reflecting a single coordinate, denoted by $\rho:\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$, e.g., $(x_{1},x_{2},\dots,x_{n})\mapsto(x_{1},x_{2},\dots,-x_{n})$, to enforce this (recall that $\mathrm{deg}(\rho)=-1$ [ref:guillemin2010differential, Ch. 3]), that is, we use $\rho\circ S$ instead of $S$. As $T$ can always be chosen to be orientation-preserving, $T$ (and equivalently $T^{-1}$) can assumed to be a stable homeomorphism [^5] The stable homeomorphism theorem—stating that all orientation-preserving homeomorphisms of $\mathbb{R}^{n}$ are stable—connects to the annulus theorem (which was a longstanding conjecture) via the work of Brown and Gluck [ref:BrownGluck] and has a rich history, with the key steps in the proof being provided by Kirby [ref:kirby1969stable] and Quinn [ref:quinn1982ends]. We point the reader to the survey of Edwards in the edited book by Gordon and Kirby on $4$-manifolds [ref:edwards1984solution]. and hence $T$ is ($C^{0}$) isotopic [^6] Indeed, as also mentioned in [ref:grune1999asymptotic], $T$ cannot (always) be a diffeomorphism on $\mathbb{R}^{n}$, for otherwise we constrain ourselves to topologically conjugate systems. to $\mathrm{id}_{\mathbb{R}^{n}}$, e.g., see [ref:kirby1969stable] and [ref:moise2013geometric, Ch. 11].
@@
(iii) First, we note that, if necessary, we can always pick the homotopy $\widetilde{H}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ from $\mathrm{id}_{\mathbb{R}^{n}}$ to $T^{-1}$ such that 00 is always mapped to 00 along the path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, e.g., for a homeomorphism $K:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $K(0)\neq 0$, consider the map $x\mapsto L(x):=K(x)-K(0)$, with $y\mapsto L^{-1}(y)=K^{-1}(y+K(0))$.
Now, recall that $0\in\mathbb{R}^{n}$ is (strongly) GAS under $\varphi^{t}(\cdot;-\nabla V)$, that is, $V(\varphi^{t}(x;-\nabla V))-V(\varphi^{0}(x;-\nabla V))\leq-\int^{t}_{0}W_{V}(\varphi^{\tau}(x;-\nabla V))\mathrm{d}\tau<0$ for all $t>0$ and any $x\neq 0$. Here, $W_{V}$ is such that $(V,W_{V})$ is a Lyapunov pair, note that in this specific case we can select $x\mapsto W_{V}(x):=\|\nabla V(x)\|_{2}^{2}$. Then, since $\widetilde{H}(s,\cdot)$ is a homeomorphism with $\widetilde{H}(s,0)=0$, we also have for any $t>0$ and $x\neq 0$ that
$V(\varphi^{t}(\widetilde{H}(s,x);-\nabla V))$ $-V(\varphi^{0}(\widetilde{H}(s,x);-\nabla V))\leq$ (3.1)
$-\int^{t}_{0}W_{V}(\varphi^{\tau}(\widetilde{H}(s,x);-\nabla V))\mathrm{d}\tau<0.$
@@
Next, define the semiflow $\widetilde{\varphi}(\cdot;s)$ through the following topological conjugacy
$\widetilde{H}(s,\cdot)\circ\widetilde{\varphi}^{t}(\cdot;s)=\varphi^{t}(\cdot;-\nabla V)\circ\widetilde{H}(s,\cdot).$

### v2

@@ 1.3   Stability
Let $\gamma\in\mathcal{K}_{\infty}$ be smooth on $(0,+\infty)$ and such that $\gamma(s)/\gamma^{\prime}(s)\geq s$ (e.g., $s\mapsto\gamma(s)=s^{1/2}$). Now consider the function $x\mapsto V_{\gamma}(x):=\gamma(\|x\|_{2})$ and construct the vector field
$\dot{x}=X_{3}(x):=\begin{cases}-\nabla V_{\gamma}(x)\quad&x\in\mathbb{R}^{n}\setminus\{0\}\\ 0\quad&\text{otherwise}\end{cases}.$ (1.5)
It seems we cannot appeal to Filippov’s framework as $X_{3}$ is not locally essentially bounded. However, as $\nabla V_{\gamma}(x)=\gamma^{\prime}(\|x\|_{2})x/\|x\|_{2}$ we can study (1.5) directly. Decompose $x\in\mathbb{R}^{n}\setminus\{0\}$ as $x=\|x\|_{2}\cdot x/\|x\|_{2}=:r\cdot u$, then, $\dot{r}=-\gamma^{\prime}(r)$ while $\dot{u}=0$. In general $\gamma^{\prime}(s)>0$ $\forall s>0$, need not be true, but suppose this is true for our choice of $\gamma$, e.g., pick $s\mapsto\gamma(s)=s^{1/2}$. Now, suppose that $1/\gamma^{\prime}(s)$ is of class $\mathcal{K}$ on $\mathbb{R}_{\geq 0}$, we can define $\Gamma(r):=\int^{r}_{0}1/\gamma^{\prime}(\rho)\mathrm{d}\rho$, which is now of class $\mathcal{K}_{\infty}$ and hence invertible on $\mathbb{R}_{\geq 0}$. Under the aforementioned assumptions, the semiflow corresponding to $X_{3}$ becomes
$(t,x)\mapsto\varphi_{3}^{t}(x):=\begin{cases}\min\{0,\Gamma^{-1}(\Gamma(\|x\|_{2})-t)\}x/\|x\|_{2}\quad&\Gamma^{-1}(t)\leq\|x\|_{2}\\ 0\quad&\mathrm{otherwise}\end{cases}.\penalty 9999\hbox{}\nobreak\hfill\quad\hbox{$\circ$}$
@@ 2   Further comments on related work
Recently, we showed that when $0\in\mathbb{R}^{n}$ is GAS under a continuous vector field $X$, and if this can be asserted using a $C^{1}$ convex Lyapunov function, then, $X$ is straight-line homotopic to $-\partial_{x}$, such that the origin remains GAS along the homotopy [ref:JongeneelSchwan2024TAC]. This is a convenient result, but not a general one.
Earlier, the following homotopy of vector fields, appeared in several works (e.g., see [ref:ryan1994brockett, p. 1603], [ref:sontag2013mathematical, Thm. 21], [ref:coron2007control, p. 291] and [ref:jongeneel2023topological, Ex. 3.4]):
@@
As for any $t\in\mathbb{R}_{>0}$, the only fixed point of $\varphi^{t}(\cdot;X)$ is 00, it follows that for $n=1$, the homotopy (2.1) preserves stability since the sign of $x\mapsto H(s,x)$ cannot flip, for otherwise, $\varphi^{s/(1-s)}(x;X)=x$ must hold for some $s\in(0,1)$ and $x\neq 0$. $\circ$
Example 2.2 (Stability preserving homotopies for linear ODEs).
Let $0\in\mathbb{R}^{n}$ be GAS under $\dot{x}=X(x):=Ax$, for $A=:TJT^{-1}$ the Jordan form decomposition of $A$. Then it follows that $\frac{1}{s}(\varphi^{s/(1-s)}(x;X)-x)=\frac{1}{s}T(e^{s/(1-s)J}-I_{n})T^{-1}x$ such that stability is preserved throughout the homotopy (2.1). $\circ$
Either way, the homotopy (2.1) does already show that there is a homotopy that does not introduce new equilibrium points, along the homotopy. Indeed, this has been generalized recently by Kvalheim to compact attractors on manifolds [ref:kvalheim2022obstructions]. On the other hand, it is known that there is no reason why stability must be preserved along such a homotopy. In the spirit of [ref:eliashberg2002introduction, Sec. 4.1], consider the following family of linear vector fields on $\mathbb{R}^{2}$:
@@
parametrizing a homotopy from $X(\cdot;0)=\partial_{x}$ to $X(\cdot;1)=-\partial_{x}$ along non-vanishing vector fields on $\mathbb{R}^{n}\setminus\{0\}$. Note, the (Hopf) indices [ref:milnor65, p. 32] of these “qualitatively opposite” vector fields are equal, i.e., $\mathrm{ind}_{0}(\partial_{x})=1^{n}=(-1)^{n}=\mathrm{ind}_{0}(-\partial_{x})$. It is this weakness of existing homotopy-invariants that we hope to overcome by studying more restrictive equivalence classes.
@@ 3   Stability preserving homotopies
Now, we will address our central research question in case $A$ is an equilibrium point. First, we consider $0\in\mathbb{R}^{n}$ being GAS under some appropriately regular vector field on $\mathbb{R}^{n}$ and eventually generalize to smooth manifolds.
@@
Our vector fields are locally essentially bounded, possibly set-valued at 00 and locally Lipschitz on $\mathbb{R}^{n}\setminus\{0\}$.
If $X$ satisfies Assumption 3.1, $\mathcal{F}[X]$ is upper semi-continuous and compact, convex valued, which allows for a smooth converse Lyapunov theory, e.g., see [ref:clarke1998asymptotic, ref:goebel2012hybrid]. We focus on Filippov’s framework, but in what follows, one may replace $\mathcal{F}[X]$ with any vector field $F$ that complies with Definition 1.3.
@@
The proof proceeds in 5 steps. First we show that the time-$t$ map $\varphi^{t}(\cdot;\mathcal{F}[X])$ can be homotoped along semiflows to a gradient flow $\varphi^{t}(\cdot;-\nabla V)$, such that along the homotopy 00 remains strongly GAS. By exploiting symmetry we show in Step (ii) that an extension of [ref:grune1999asymptotic] allows for showing that for any $\gamma\in\mathcal{K}_{\infty}$ there is a $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ (not just $T\in\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$) such that $V(T^{-1}(x))=\gamma(\|x\|_{2})$. Then, it follows that $T$ is homotopic to $\mathrm{id}_{\mathbb{R}^{n}}$ along a path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$. In Step (iii), we use this map $T$ to homotope $\varphi^{t}(\cdot;-\nabla V)$ to a semiflow, denoted $\widetilde{\varphi}$, that has $V\circ T^{-1}$ as its Lyapunov function. Again, 00 remains strongly GAS along the path. Next, we show in Step (iv) that $\widetilde{\varphi}^{t}$ can be homotoped to $\varphi^{t}(\cdot;-\partial_{x})$. We conclude in Step (v).
(i) Since 00 is strongly GAS under $\dot{x}\in\mathcal{F}[X](x)$ and $X$ satisfies Assumption 3.1, there is a $C^{\infty}$ Lyapunov pair $(V,W)$, that certifies stability of $0\in\mathbb{R}^{n}$ [ref:clarke1998asymptotic, Thm. 1.3], under any Filippov solution. Then, construct the homotopy $H:[0,1]\times\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}^{n}\setminus\{0\}$ through non-vanishing vector fields, defined by $(s,x)\mapsto H(s,x):=(1-s)\mathcal{F}[X](x)-s\nabla V(x)$. As for any $s\in[0,1]$, $(1-s)\mathcal{F}[X]-s\nabla V$ satisfies Assumption 3.1, plus $\langle\nabla V(x),H(s,x)\rangle\leq-(1-s)W(x)-s\|\nabla V(x)\|_{2}^{2}$ $\forall\,x\in\mathbb{R}^{n}\setminus\{0\}$, we find that 00 is strongly GAS, in the sense of Filippov, under $\dot{x}\in(1-s)\mathcal{F}[X](x)-s\nabla V(x)$, for any fixed $s\in[0,1]$. To construct a homotopy on the level of semiflows, we follow Hartman [hartman2002ordinary, p. 93] and consider the following extended system:
@@
As $V$ has compact sublevel sets (see Definition 1.3), any Filippov solution to $\Sigma_{H}$ is defined for all $t\geq 0$, e.g., see [ref:khalil2002nonlinear, Thm. 3.3]. Now since $(s,x)\mapsto H(s,x)$ is locally Lipschitz away from 00, we can appeal to the Picard-Lindelöf theorem [ref:teschl2012ordinary, Thm. 2.2] and hence, by strong asymptotic stability [ref:clarke1998asymptotic, Def. 2.1] cf. Example 1.4, any solution to $\Sigma_{H}$ is also unique, which allows us to define the time-$t$ map $(s,x)\mapsto\varphi^{t}((s,x);\Sigma_{H})$ for any $t\geq 0$. In its turn, by considering the path $[0,1]\ni s\mapsto\varphi^{t}((s,\cdot);\Sigma_{H})$, we see that this time-$t$ map $\varphi^{t}$ defines a homotopy from the semiflow under $\mathcal{F}[X]$ to the (semi)flow under $-\nabla V$. Specifically, we have that $\varphi^{t}((0,x);\Sigma_{H})=(0,\varphi^{t}(x;\mathcal{F}[X]))$, $\varphi^{t}((1,x);\Sigma_{H})=(1,\varphi^{t}(x;-\nabla V))$ and $\varphi^{t}((s,0);\Sigma_{H})=(s,0)$ $\forall s\in[0,1]$. Now, for any fixed $s\in[0,1]$, overload notation and define the time-$t$ map $\varphi^{t}(\cdot;(s,\Sigma_{H})):=\pi_{2:n+1}\circ\varphi^{t}((s,\cdot);\Sigma_{H})$, for $\pi_{2:n+1}$ the projection on the last $n$ coordinates. It follows that this map checks out as a semiflow, since $\pi_{2:n+1}$ is continuous and for any $x\in\mathbb{R}^{n}$ we have that
@@
(ii) Grüne, Wirth and Sontag showed that for our Lyapunov function $V$, we can find a $T\in\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$, with $T(0)=0$, such that $V(T^{-1}(x))=\gamma(\|x\|_{2})$, for some $\gamma\in\mathcal{K}_{\infty}$, smooth on $(0,+\infty)$ [ref:grune1999asymptotic, Prop. 1]. We recall their construction $T$. Define $(t,x)\mapsto\psi(t,x)$ to be the flow with respect to
@@
It follows that, on the domain of $\psi$, $V(\psi(t,x))=V(x)+t$. Now fix some $c>0$, then define the map $\pi_{c}:\mathbb{R}^{n}\setminus\{0\}\to V^{-1}(c)$ by $x\mapsto\pi_{c}(x):=\psi(c-V(x),x)$, that is, starting from $x\in\mathbb{R}^{n}\setminus\{0\}$ you flow—either backward or forward—along $\psi$ until you hit $V^{-1}(c)$. Then, due to initial work by Wilson [ref:wilson1967structure], we know that $V^{-1}(c)\simeq_{h}\mathbb{S}^{n-1}$, however, now—that is, after [ref:grune1999asymptotic] was published, the resolution of the Poincaré conjecture [^3] Specifically, Perelman provided the final step $(\mathbb{S}^{3})$ in proving the generalized Poincaré conjecture in $\mathsf{Top}$. For some historical comments, see [ref:stillwell2012poincare]. implies there must be a homeomorphism $S:V^{-1}(c)\to\mathbb{S}^{n-1}$, for any $n\geq 1$. Next, define the map $Q:=S\circ\pi_{c}:\mathbb{R}^{n}\setminus\{0\}\to\mathbb{S}^{n-1}$ and eventually the map $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by
@@
$x\mapsto T(x):=\begin{cases}\gamma^{-1}(V(x))Q(x)\quad&x\in\mathbb{R}^{n}\setminus\{0\}\\ 0\quad&\text{otherwise},\end{cases}$
@@
Now, it turns out that the particular choice of the homeomorphism $S:V^{-1}(c)\to\mathbb{S}^{n-1}$ is irrelevant for the construction of $T$ as in [ref:grune1999asymptotic, Prop. 1], as such, we select $S$ to enforce $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$. If $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, we are done, if not, we can compose $S$ with a map reflecting a single coordinate, denoted by $\rho:\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$, e.g., $(x_{1},x_{2},\dots,x_{n})\mapsto(x_{1},x_{2},\dots,-x_{n})$, to change the orientation of $T$ (recall that $\mathrm{deg}(\rho)=-1$ [ref:guillemin2010differential, Ch. 3]), that is, we use $\rho\circ S$ instead of $S$. Thus, as $T$ can always be chosen to be orientation-preserving, $T$ (and equivalently $T^{-1}$) can assumed to be a stable homeomorphism [^4] The stable homeomorphism theorem—stating that all orientation-preserving homeomorphisms of $\mathbb{R}^{n}$ are stable—connects to the annulus theorem (which was a longstanding conjecture) via the work of Brown and Gluck [ref:BrownGluck] and has a rich history, with the key steps in the proof being provided by Kirby [ref:kirby1969stable] and Quinn [ref:quinn1982ends]. We point the reader to the survey of Edwards in the edited book by Gordon and Kirby on $4$-manifolds [ref:edwards1984solution]. and hence $T$ is isotopic [^5] Indeed, as also mentioned in [ref:grune1999asymptotic], $T$ cannot (always) be a diffeomorphism on $\mathbb{R}^{n}$, for otherwise we constrain ourselves to smoothly conjugate systems. to $\mathrm{id}_{\mathbb{R}^{n}}$, e.g., see [ref:kirby1969stable] and [ref:moise2013geometric, Ch. 11].
@@
(iii) First, we can always pick the homotopy $\widetilde{H}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ from $\mathrm{id}_{\mathbb{R}^{n}}$ to $T^{-1}$ such that 00 is mapped to 00 along the path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, e.g., for a homeomorphism $K:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $K(0)\neq 0$, consider the map $x\mapsto L(x):=K(x)-K(0)$, with $y\mapsto L^{-1}(y)=K^{-1}(y+K(0))$. Now recall that $0\in\mathbb{R}^{n}$ is GAS under $\varphi^{t}(\cdot;-\nabla V)$, i.e., $V(\varphi^{t}(x;-\nabla V))-V(x)\leq-\int^{t}_{0}W_{V}(\varphi^{\tau}(x;-\nabla V))\mathrm{d}\tau$ $\forall(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$, with $x\mapsto W_{V}(x):=\|\nabla V(x)\|_{2}^{2}$. Since $\widetilde{H}(s,\cdot)$ is a homeomorphism with $\widetilde{H}(s,0)=0$, we also have
$V(\varphi^{t}(\widetilde{H}(s,x);-\nabla V))-V(\widetilde{H}(s,x))\leq-\int^{t}_{0}W_{V}(\varphi^{\tau}(\widetilde{H}(s,x);-\nabla V))\mathrm{d}\tau,$
@@
for any $(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$. Next, define the semiflow $\widetilde{\varphi}(\cdot;s)$ through
$\widetilde{H}(s,\cdot)\circ\widetilde{\varphi}^{t}(\cdot;s)=\varphi^{t}(\cdot;-\nabla V)\circ\widetilde{H}(s,\cdot).$ (3.1)